OpenStudy (anonymous):

1.) x+4/6=6 2.) 9x-4=11x-8+x 3.) You buy 2 erasers and 18 for 0.8 cents each. Your friend buys 7 erasers and 15 pencils. You both end up paying the same amount. How much are each eraser?

OpenStudy (texaschic101):

is the first one : x + (4/6) = 6 or (x + 4)/6 = 6 ??

OpenStudy (texaschic101):

2. 9x - 4 = 11x - 8 + x -- combine like terms 9x - 4 = 12x - 8 -- add 8 to both sides 9x - 4 + 8 = 12x -- subtract 9x from both sides -4 + 8 = 12x - 9x -- combine like terms 4 = 3x -- divide by 3 4/3 = x 3. you need to make this question a little more clearer. I do not understand it.